A pole 24 feet high has a shadow 8 feet long. A nearby pole is 72 feet tall, how long is its shadow?
Do I set this up as a ratio? 8/24...x/72? cross mulitply?

Answer by stanbon(75887)

A pole 24 feet high has a shadow 8 feet long. A nearby pole is 72 feet tall, how long is its shadow?
Do I set this up as a ratio? 8/24...x/72? cross mulitply?
---
Yes, you do.
Let the shadow have length "x".
Proportion:
x/72 = 8/24
---
x = 72(8/24)
x = 24 ft.
=================
Cheers,
Stan H.